OpenStudy (anonymous):

How do I find the equation of a line having the given slope of m=5/6 and the given points (8,-8). I am completely confused and any help would be greatly appreciated.

OpenStudy (anonymous):

Think about what does the slope means. It means that equation has to start with something like y=mx +some number Now you have the point given, (8,-8) that means that at the point x=8, y=-8 put these values in the equation and count that +some number. Yell if something is not clear

OpenStudy (anonymous):

that is one clear way. also don't forget the all mighty 'point-slope' formula which says that a line through \[(x_1,y_1)\] with slope \[m\] has the equation \[y-y_1=m(x-x_1)\] where of course \[x_1, y_1\] are numbers and x and y are the variables in your anwer. so you can write \[y-(-8)=\frac{5}{6}(x-8)\] and solve for y